A heating coil of resistance 20 ohms is connected to a 220 V supply for 2 mins to boi a quantity of water in a vessel of heat capacity 100 JK-1. The initial temperature of the water is 400C and its specific heat capacity is 4.2 x 103Jkg-1K-1. calculate the mass of water in the vessel.
1
Expert's answer
2020-06-03T12:10:55-0400

Calculate the energy required to boil the mass m of water. The electrical energy "P\tau" from the heater will heat up the vessel (heat required is "C\Delta t") and the water (heat required "cm\Delta t"):


"Q=cm\Delta t+C\Delta t=P\tau,\\
cm\Delta t+C\Delta t=\frac{V^2}{R}\tau,\\\space\\
m=\frac{1}{c}\bigg(\frac{V^2\tau}{R\Delta t}-C\bigg)=\\\space\\
=\frac{1}{4200}\bigg(\frac{220^2\cdot2\cdot60}{20(100-40)}-100\bigg)=1.13\text{ kg}."